OpenStudy (anonymous):

Simply the expression: (2^2)(2^5)

OpenStudy (kira_yamato):

2^7 = 128

OpenStudy (anonymous):

2^2 means 2*2 and 2^5 means 2*2*2*2*2 so =(2*2)(2*2*2*2*2) =(4)(32) =128

OpenStudy (anonymous):

2^7 (2^2+5) = 2^7 = 128